# Classical and quantum thermodynamics in a non–equilibrium regime: Application to Stirling engine

### v1

@@ II.1 Laws of thermodynamics as applied to work
which corresponds to the principle of maximum entropy, where $\Lambda_{\rm A}^{\rm qst}[U_{\rm A}^{\rm qst},{X}_{\rm A}^{\rm qst}]$ is the entropic potential. Because both natural variables are extensive, this potential is fundamental and we call it Massieu entropy (M-entropy). Using the TDL tranformation Eq. (46), we also have Planck entropy (P-entropy) defined by $\Gamma_{\rm A}^{\rm qst}[U_{\rm A}^{\rm qst},\tilde{x}^{\rm qst}]$ from $\tilde{Q}_{\rm A}^{int}(t)$ (see Appendix A). Similar to the entropy obtained from the partition function in statistical physics, this entropy is a function of the extensive variable $\tilde{x}^{\rm qst}$.
Here, as the conjugate variable for $\beta(t)$, we chose the internal energy $U_{\rm A}^{\rm qst}$, which is considered a fundamental variable in thermodynamics, but enthalpy $H_{\rm A}^{\rm qst}$ could also be chosen.Koyanagi and Tanimura (2024a) In such case, there are also two entropies, one involving both extensive variables (Clasius entropy or C-entropy) and one for intensive variable with respect to external forces (Boltzmann entropy or B-entropy). The B-entropy and M-entopy are related to the others by the Legendre transformations between $U_{\rm A}^{\rm qst}$ and $H_{\rm A}^{\rm qst}$ and the values do not change, although the former inculdes one intensive variable whereas the later includes extensive variables only. However, there is no explicit relationship between C-entropy and M-entropy, because the Legendre transformation between two entropy in the $U_{\rm A}^{\rm qst}$ representation and the $H_{\rm A}^{\rm qst}$ representation are different. (See Appendix D in Ref. Koyanagi and Tanimura, 2024a). In any case, the Massieu–Planck potentials are entropic potentials.Guggenheim (1986)
The total differential form of the Helmholtz–Gibbs potentials can be derived from the Massieu–Planck potentials using the definitions $F^{\rm qst}_{\rm A}(t)=-\Phi_{\rm A}^{\rm qst}(t)/\beta^{\rm qst}(t)$ and $G^{\rm qst}_{\rm A}(t)=-\Xi_{\rm A}^{\rm qst}(t)/\beta^{\rm qst}(t)$. The results are summarized and presented in Table 6 in Appendix A.
II.2 Thermodynamic potentials in a non–equilibrium regime
II.2.1 The DL non–equilibrium minimum work principle
On the basis of Eq. (8), we can define thermodynamic potentials applicable to the non–equilibrium regime. Consider a non–equilibrium state $A$ and equilibrium state $n$. The DL non–equilibrium-to-equilibrium minimum work principle for $A\rightarrow n$ is expressed as (see Appendix B)
@@ II.2.1 The DL non–equilibrium minimum work principle
where $\Phi_{\rm A}^{\rm neq}(t)$ is the non–equilibrium Massieu potential. This inequality indicates that the path from non–equilibrium to equilibrium has a lower bound on work. In engineering, the effective energy from non-equilibrium to equilibrium is referred to as exergy.Rant (1956); Rosen and Dincer (2001) The non–equilibrium thermodynamic potential we introduced here can be regarded as a generalization of it.
The difference between $(\tilde{W}_{\rm A}^{ext})_{{A}\rightarrow n}$ and $-(\Delta\Phi_{A}^{\rm neq})_{{A}\rightarrow n}$ corresponds to the DL heat generated as loss work, denoted by $-(Q_{\rm A}^{\rm wst})_{{A}\rightarrow n}$, which we call DL waste heat:
$-(\tilde{Q}_{\rm A}^{\rm wst})_{{A}\rightarrow n}=(\tilde{W}_{\rm A}^{ext})_{{A}\rightarrow n}+(\Delta\Phi_{\rm A}^{\rm neq})_{{A}\rightarrow n}\geq 0.$ (11)
Because the DL waste heat between two non–equilibrium states $A$ and $B$ is evaluated as $(\tilde{Q}_{\rm A}^{\rm wst})_{{A}\rightarrow{B}}=(\tilde{Q}_{\rm A}^{\rm wst})_{{A}\rightarrow n}-(\tilde{Q}_{\rm A}^{\rm wst})_{B\rightarrow n}$, we the waste heat for any non–equilibrium process at times $t$ and $t+dt$ (where $dt$ is an infinitesimal time) in time-derivative form:
$-\frac{d\tilde{Q}_{\rm A}^{\rm wst}(t)}{dt}=\frac{d\tilde{W}_{\rm A}^{ext}(t)}{dt}+\frac{d\Phi_{\rm A}^{\rm neq}(t)}{dt}.$ (12)
As shown in Appendix B, the waste heat is nonpositive, i.e., $d\tilde{Q}_{\rm A}^{\rm wst}(t)/dt\leq 0$. Thus, we have the inequality $(\tilde{W}_{\rm A}^{ext})_{A\rightarrow B}^{\rm min}\geq-(\Delta\Phi_{\rm A}^{\rm neq})_{\rm A\rightarrow B}$, or
@@
which we call the DL non–equilibrium minimum work principle. While $\tilde{W}_{\rm A}^{ext}(t)$ and $\tilde{Q}_{\rm A}^{\rm wst}(t)$ are not state variables because they depend on a path, $\Phi_{\rm A}^{\rm neq}(t)$ is a state variable defined by the non–equilibrium-to-equilibrium minimum work path. When state $B$ is on the non–equilibrium-to-equilibrium minimum work path from $A$ to $n$, equality holds in (13).
Using Eqs. (2) and (12), we obtain the time-derivative form of the non–equilibrium Massieu potential as
$\frac{d\Phi_{\rm A}^{\rm neq}(t)}{dt}=-U_{\rm A}^{\rm neq}\frac{d\beta(t)}{dt}-\tilde{x}(t)\frac{d{X}_{\rm A}^{\rm neq}(t)}{dt}-\frac{d\tilde{Q}_{\rm A}^{\rm wst}(t)}{dt}.$ (14)
The convexity of DL potentials is discussed on the basis of the SB model in Appendix C. Thus, we can regard these as thermodynamic potentials.
Using Eqs. (6), we have the non–equilibrium Planck potential expressed as
$\frac{d\Xi_{\rm A}^{\rm neq}(t)}{dt}=-U_{\rm A}^{\rm neq}\frac{d\beta(t)}{dt}+{X}_{\rm A}^{\rm neq}(t)\frac{d\tilde{x}(t)}{dt}-\frac{d\tilde{Q}_{\rm A}^{\rm wst}(t)}{dt}.$ (15)
The M-entropy $\Lambda_{\rm A}^{\rm neq}(t)$ and P-entropy $\Gamma_{\rm A}^{\rm neq}(t)$ can be evaluated from the TDL transformations (6) and  (7). The non–equilibrium potentials satisfy the Legendre transformations:
@@
The time-derivative forms of the DL non–equilibrium entropic potentials are summarized in Table 1. Note that while the entropic potentials, the intensive variables and non–equilbrium extensive variables are state variables, the time–deribative of them including waste heat are not state variables.
As shown in Appendix D, the non–equilibrium Massieu potential is always smaller than the quasi-static Massieu potential. This indicates that the non–equilibrium Massieu potential is minimum when the state is equilibrium. In the quasi-static case, Table  1 is reduced to Table 5 (Appendix A). Thus, we can regard these potentials as extensions of the thermodynamic potentials to a non–equilibrium regime.
From Eqs. (12) and (17), the waste heat current is expressed using the non–equilibrium M-entropy as
$\frac{d\tilde{Q}_{\rm A}^{\rm wst}(t)}{dt}=\frac{d\tilde{Q}_{\rm A}^{ext}(t)}{dt}-\frac{d\Lambda_{\rm A}^{\rm neq}(t)}{dt}.$ (18)
Table 1: Time-derivative forms of the non–equilibrium (neq) Massieu–Planck potentials as functions of the intensive variables $\beta(t)$ and $\tilde{x}(t)$ and the extensive variables $U_{\rm A}^{\rm neq}(t)$ and $X_{\rm A}^{\rm neq}(t)$. Of the DL entropies, the commonly used one, which we call the Massieu entropy (M-entropy), involves only extensive variables and is denoted by $\Lambda_{\rm A}^{\rm neq}[U_{\rm A}^{\rm neq}(t),{X}_{\rm A}^{\rm neq}(t)]$, whereas the less widely used one, which we call the Planck entropy (P-entropy), is denoted by $\Gamma_{\rm A}^{\rm qst}[U_{\rm A}^{\rm qst},\tilde{x}]$. Each potential is related to the others by the Legendre transformations shown in the final column.
@@
Massieu $\frac{d}{dt}\Phi_{\rm A}^{\rm neq}=-U_{\rm A}^{\rm neq}\frac{d}{dt}\beta-\tilde{x}\frac{d}{dt}{X}_{\rm A}^{\rm neq}-\frac{d}{dt}\tilde{Q}^{\rm wst}_{\rm A}$ $\beta,{X}_{\rm A}^{\rm neq}$cdots
Planck $\frac{d}{dt}\Xi^{\rm neq}_{\rm A}=-U_{\rm A}^{\rm neq}\frac{d}{dt}\beta+{X}_{\rm A}^{\rm neq}\frac{d}{dt}\tilde{x}-\frac{d}{dt}\tilde{Q}^{\rm wst}_{\rm A}$ $\beta,\tilde{x}$ $\Xi_{\rm A}^{\rm neq}=\Phi_{\rm A}^{\rm neq}+\tilde{x}{X}_{\rm A}^{\rm neq}$
M-Entropy $\frac{d}{dt}\Lambda_{\rm A}^{\rm neq}=\beta\frac{d}{dt}U_{\rm A}^{\rm neq}-\tilde{x}\frac{d}{dt}X_{\rm A}^{\rm neq}-\frac{d}{dt}\tilde{Q}_{\rm A}^{\rm wst}$ $U_{\rm A}^{\rm neq},X_{\rm A}^{\rm neq}$ $\Lambda_{\rm A}^{\rm neq}=\Phi_{\rm A}^{\rm neq}+\beta U_{\rm A}^{\rm neq}$
P-Entropy $\frac{d}{dt}\Gamma_{\rm A}^{\rm neq}=\beta\frac{d}{dt}U_{\rm A}^{\rm neq}+X_{\rm A}^{\rm neq}\frac{d}{dt}\tilde{x}-\frac{d}{dt}\tilde{Q}_{\rm A}^{\rm wst}$ $U_{\rm A}^{\rm neq},\tilde{x}$ $\Gamma_{\rm A}^{\rm neq}=\Lambda_{\rm A}^{\rm neq}+\tilde{x}{X}_{\rm A}^{\rm neq}$
II.2.2 Nonequilibrium Helmholtz–Gibbs potentials
Table 2: Time-derivative forms of the non–equilibrium (neq) Helmholtz–Gibbs potentials as functions of the intensive variables $T(t)$ and $x(t)$ and the extensive variables $S_{\rm A}^{\rm neq}(t)$ and $X_{\rm A}^{\rm neq}(t)$, which are interrelated through the Legendre transformations shown in the final column. Because heat is always lost in non–equilibrium processes, the waste heat $Q^{\rm wst}(t)$ appears in the equations.
@@ II.2.2 Nonequilibrium Helmholtz–Gibbs potentials
Helmholtz Energy $\frac{d}{dt}F^{\rm neq}_{\rm A}=-S^{\rm neq}_{\rm A}\frac{d}{dt}T+x\frac{d}{dt}X_{\rm A}^{\rm neq}+\frac{d}{dt}Q^{\rm wst}$ $T,x$
Gibbs Energy $\frac{d}{dt}G^{\rm neq}_{\rm A}=-S^{\rm neq}_{\rm A}\frac{d}{dt}T-X_{\rm A}^{\rm neq}\frac{d}{dt}x+\frac{d}{dt}Q^{\rm wst}$ $T,X_{\rm A}^{\rm neq}$ $G^{\rm neq}_{\rm A}=F^{\rm neq}_{\rm A}-xX_{\rm A}^{\rm neq}$
Internal Energy $\frac{d}{dt}U^{\rm neq}_{\rm A}=T\frac{d}{dt}S^{\rm neq}_{\rm A}+x\frac{d}{dt}X_{\rm A}^{\rm neq}+\frac{d}{dt}Q^{\rm wst}$ $S_{\rm A}^{\rm neq},x$ $U^{\rm neq}_{\rm A}=F^{\rm neq}_{\rm A}+TS^{\rm neq}_{\rm A}$
Enthalpy $\frac{d}{dt}H^{\rm neq}_{\rm A}=T\frac{d}{dt}S^{\rm neq}_{\rm A}-X_{\rm A}^{\rm neq}\frac{d}{dt}x+\frac{d}{dt}Q^{\rm wst}$ $S_{\rm A}^{\rm neq},X_{\rm A}^{\rm neq}$ $H^{\rm neq}_{\rm A}=U^{\rm neq}_{\rm A}-xX_{\rm A}^{\rm neq}$
We introduce non–equilibrium Helmholtz and Gibbs energies defined as $F^{\rm neq}_{\rm A}(t)=-\Phi_{\rm A}^{\rm neq}(t)/\beta(t)$ and $G^{\rm neq}_{\rm A}(t)=-\Xi_{\rm A}^{\rm neq}(t)/\beta(t)$. Because $d\beta(t)/dt=-(1/k_{\rm B}T^{2}(t))dT(t)/dt$, we obtain the time-derivative forms of these from Eqs. (14) and (15) as
@@
$\frac{dG_{\rm A}^{\rm neq}(t)}{dt}=-S_{\rm A}^{\rm neq}(t)\frac{dT(t)}{dt}+X_{\rm A}^{\rm neq}(t)\frac{dx(t)}{dt}+\frac{dQ_{\rm A}^{\rm wst}(t)}{dt},$ (20)
where $S_{\rm A}^{\rm neq}(t)=k_{\rm B}\Lambda_{\rm A}^{\rm neq}(t)$ is a non–equilibrium entropy and we have introduced the waste heat current as
$\frac{dQ_{\rm A}^{\rm wst}(t)}{dt}=\frac{1}{\beta(t)}\frac{d\tilde{Q}_{\rm A}^{\rm wst}(t)}{dt}.$ (21)
From Eq. (16), we obtain the TDL transformation between the non–equilibrium Gibbs and Helmholtz potentials as
@@
Solving Eq. (17) for $U_{\rm A}^{\rm neq}(t)$, we obtain the following TDL transformations:
@@
From the above, we obtain the time-derivative expressions for the enthalpy and internal energy. The results are summarized in Table 2. These results reduce to those in Table 6 (Appendix A) in the quasi-static case. Thus, we can regard these potentials as extensions of the thermodynamic potentials to a non–equilibrium regime.
@@
is the extensive heat current. From Eqs. (18) and (25), we can evaluate the DL waste heat current as
$-\frac{d\tilde{Q}_{\rm A}^{\rm wst}(t)}{dt}=\frac{1}{k_{\rm B}}\left\{\frac{dS_{\rm A}^{\rm neq}(t)}{dt}-\frac{1}{T(t)}\frac{dQ_{\rm A}^{ext}(t)}{dt}\right\}.$ (26)
In the isothermal case, by integrating both sides between the two equilibrium states over time $t$, we obtain
$-\tilde{Q}_{\rm A}^{\rm wst}=\frac{1}{k_{\rm B}}\left\{\Delta S_{\rm A}^{\rm qst}-\frac{Q_{\rm A}^{ext}}{T}\right\},$ (27)
which indicates that the DL waste heat is equivalent to the entropy production evaluated for a heat engine.Sakamoto and Tanimura (2020); Koyanagi and Tanimura (2022a)
The non–equilibrium Gibbs energy satisfies
@@
$H_{\rm A}(t)=-T^{2}(t)\frac{\partial}{\partial T(t)}\left(\frac{G^{\rm neq}_{\rm A}(t)}{T(t)}\right)_{x(t),\tilde{Q}^{\rm wst}}$ (30)
Equation (30) is an extension of the Gibbs–Helmholtz relation to a non–equilibrium regime.

### v2

@@ II.1 Laws of thermodynamics applied to work
which corresponds to the principle of maximum entropy, where $\Lambda_{\rm A}^{\rm qst}[U_{\rm A}^{\rm qst},{X}_{\rm A}^{\rm qst}]$ is the entropic potential. Because both natural variables are extensive, this potential is fundamental and we call it the Massieu entropy (M-entropy). Using the TDL transformation given by Eq. (46), we also obtain the Planck entropy (P-entropy) defined by $\Gamma_{\rm A}^{\rm qst}[U_{\rm A}^{\rm qst},\tilde{x}^{\rm qst}]$ from $\tilde{Q}_{\rm A}^{int}(t)$ (see Appendix A). Similar to the entropy obtained from the partition function in statistical physics, this entropy is a function of the extensive variable $\tilde{x}^{\rm qst}$.
Here, as the conjugate variable for $\beta(t)$, we chose the internal energy $U_{\rm A}^{\rm qst}$, which is considered a fundamental variable in thermodynamics. As demonstrated in our previous paper,Koyanagi and Tanimura (2024b) enthalpy $H_{\rm A}^{\rm qst}$ could also be chosen as the conjugate variable for $\beta(t)$. In such a case, there are also two entropies, one involving both extensive variables (Clausius entropy or C-entropy) and one intensive variable with respect to external forces (Boltzmann entropy or B-entropy). The B-entropy and M-entropy values are related to the others by Legendre transformations between $U_{\rm A}^{\rm qst}$ and $H_{\rm A}^{\rm qst}$ and the values do not change, although the former includes one intensive variable, whereas the later includes extensive variables only. However, there is no explicit relationship between C-entropy and M-entropy, because the Legendre transformations between two entropies in the $U_{\rm A}^{\rm qst}$ representation and the $H_{\rm A}^{\rm qst}$ representation are different. (See Appendix D in Ref. Koyanagi and Tanimura, 2024b). In both cases, the Massieu and Planck potentials are the same entropic potentials.Guggenheim (1986)
The total differential form of the Helmholtz–Gibbs potentials can be derived from the Massieu–Planck potentials using the definitions $F^{\rm qst}_{\rm A}(t)=-\Phi_{\rm A}^{\rm qst}(t)/\beta^{\rm qst}(t)$ and $G^{\rm qst}_{\rm A}(t)=-\Xi_{\rm A}^{\rm qst}(t)/\beta^{\rm qst}(t)$. The results are summarized in Table 6 in Appendix A.
II.2 Thermodynamic potentials in a non-equilibrium regime
II.2.1 The DL non-equilibrium minimum work principle
On the basis of Eq. (8), we define thermodynamic potentials applicable to the non-equilibrium regime. Consider a non-equilibrium state $A$ and equilibrium state $n$. The DL non-equilibrium-to-equilibrium minimum work principle for $A\rightarrow n$ is expressed as (see Appendix B)
@@ II.2.1 The DL non-equilibrium minimum work principle
where $\Phi_{\rm A}^{\rm neq}(t)$ is the non-equilibrium Massieu potential. This inequality indicates that the path from non-equilibrium to equilibrium has a lower bound on work. In engineering, the effective energy from non-equilibrium to equilibrium is referred to as exergy.Z The non-equilibrium thermodynamic potential introduced here can be regarded as a generalization of it.
The difference between $(\tilde{W}_{\rm A}^{ext})_{{A}\rightarrow n}$ and $-(\Delta\Phi_{A}^{\rm neq})_{{A}\rightarrow n}$ corresponds to the entropy production defined as
$(\Sigma_{\rm A})_{{A}\rightarrow n}=(\tilde{W}_{\rm A}^{ext})_{{A}\rightarrow n}+(\Delta\Phi_{\rm A}^{\rm neq})_{{A}\rightarrow n}\geq 0.$ (11)
For two non-equilibrium states $A$ and $B$, we have $(\Sigma_{\rm A})_{{A}\rightarrow B}$ = $(\Sigma_{\rm A})_{{A}\rightarrow n}-(\Sigma_{\rm A})_{{B}\rightarrow n}$. Thus, the entropy production rate for any non-equilibrium process at times $t$ and $t+dt$ (where $dt$ is an infinitesimal time) is expressed as
$\frac{d\Sigma_{\rm A}(t)}{dt}=\frac{d\tilde{W}_{\rm A}^{ext}(t)}{dt}+\frac{d\Phi_{\rm A}^{\rm neq}(t)}{dt}.$ (12)
As shown in Appendix B, the entropy production rate is alway positive, i.e., $d\Sigma_{\rm A}(t)/dt\geq 0$. Thus, we have the inequality $(\tilde{W}_{\rm A}^{ext})_{A\rightarrow B}^{\rm min}\geq-(\Delta\Phi_{\rm A}^{\rm neq})_{\rm A\rightarrow B}$, or
@@
which we call the DL non-equilibrium minimum work principle. While $\tilde{W}_{\rm A}^{ext}(t)$ and $\Sigma_{\rm A}(t)$ are not state variables because they depend on a path, $\Phi_{\rm A}^{\rm neq}(t)$ is a state variable defined by the non-equilibrium-to-equilibrium minimum work path. When state $B$ is on the non-equilibrium-to-equilibrium minimum work path from $A$ to $n$, equality in Eq. (13) holds.
Using Eqs. (2) and (12), we obtain the time derivative form of the non-equilibrium Massieu potential as
$\frac{d\Phi_{\rm A}^{\rm neq}(t)}{dt}=-U_{\rm A}^{\rm neq}\frac{d\beta(t)}{dt}-\tilde{x}(t)\frac{d{X}_{\rm A}^{\rm neq}(t)}{dt}+\frac{d\Sigma_{\rm A}(t)}{dt}.$ (14)
The convexity of DL potentials is discussed on the basis of the SB model in Appendix C. Thus, these can be regarded as thermodynamic potentials.
Using Eqs. (6), we have the non-equilibrium Planck potential expressed as
$\frac{d\Xi_{\rm A}^{\rm neq}(t)}{dt}=-U_{\rm A}^{\rm neq}\frac{d\beta(t)}{dt}+{X}_{\rm A}^{\rm neq}(t)\frac{d\tilde{x}(t)}{dt}+\frac{d\Sigma_{\rm A}(t)}{dt}.$ (15)
The M-entropy $\Lambda_{\rm A}^{\rm neq}(t)$ and P-entropy $\Gamma_{\rm A}^{\rm neq}(t)$ can be evaluated from the TDL transformations (6) and  (7). The non-equilibrium potentials satisfy the following Legendre transformations:
@@
The time derivative forms of the DL non-equilibrium entropic potentials are summarized in Table 1. Note that the entropic potentials, the intensive variables, and the non-equilibrium extensive variables are state variables, whereas their time derivatives are not state variables.
As shown in Appendix D, the non-equilibrium Massieu potential is always smaller than the quasi-static Massieu potential. This indicates that the non-equilibrium Massieu potential is minimum when the state is equilibrium. In the quasi-static case, Table  1 is reduced to Table 5 (Appendix A). Thus, we can regard these potentials as extensions of the thermodynamic potentials to a non-equilibrium regime.
From Eqs. (12) and (17), the entropy production rate is expressed using the non-equilibrium M-entropy as
$\frac{d\Sigma_{\rm A}(t)}{dt}=\frac{d\Lambda_{\rm A}^{\rm neq}(t)}{dt}-\frac{d\tilde{Q}_{\rm A}^{ext}(t)}{dt}.$ (18)
Table 1: Time derivative forms of the non-equilibrium (neq) Massieu–Planck potentials as functions of intensive variables $\beta(t)$ and $\tilde{x}(t)$ and extensive variables $U_{\rm A}^{\rm neq}(t)$ and $X_{\rm A}^{\rm neq}(t)$. Of the DL entropies, the commonly used one, which we call the Massieu entropy (M-entropy), involves only extensive variables and is denoted by $\Lambda_{\rm A}^{\rm neq}[U_{\rm A}^{\rm neq}(t),{X}_{\rm A}^{\rm neq}(t)]$, whereas the less widely used one, which we call the Planck entropy (P-entropy), is denoted by $\Gamma_{\rm A}^{\rm qst}[U_{\rm A}^{\rm qst},\tilde{x}]$. Because heat is always lost in non-equilibrium processes, the entropy production rate $d\Sigma_{\rm A}/dt$ appears in the equations. Each potential is related to others via Legendre transformations shown in the final column.
@@
Massieu $\frac{d}{dt}\Phi_{\rm A}^{\rm neq}=-U_{\rm A}^{\rm neq}\frac{d}{dt}\beta-\tilde{x}\frac{d}{dt}{X}_{\rm A}^{\rm neq}+\frac{d}{dt}\Sigma_{\rm A}$ $\beta,{X}_{\rm A}^{\rm neq}$cdots
Planck $\frac{d}{dt}\Xi^{\rm neq}_{\rm A}=-U_{\rm A}^{\rm neq}\frac{d}{dt}\beta+{X}_{\rm A}^{\rm neq}\frac{d}{dt}\tilde{x}+\frac{d}{dt}\Sigma_{\rm A}$ $\beta,\tilde{x}$ $\Xi_{\rm A}^{\rm neq}=\Phi_{\rm A}^{\rm neq}+\tilde{x}{X}_{\rm A}^{\rm neq}$
M-Entropy $\frac{d}{dt}\Lambda_{\rm A}^{\rm neq}=\beta\frac{d}{dt}U_{\rm A}^{\rm neq}-\tilde{x}\frac{d}{dt}X_{\rm A}^{\rm neq}+\frac{d}{dt}\Sigma_{\rm A}$ $U_{\rm A}^{\rm neq},X_{\rm A}^{\rm neq}$ $\Lambda_{\rm A}^{\rm neq}=\Phi_{\rm A}^{\rm neq}+\beta U_{\rm A}^{\rm neq}$
P-Entropy $\frac{d}{dt}\Gamma_{\rm A}^{\rm neq}=\beta\frac{d}{dt}U_{\rm A}^{\rm neq}+X_{\rm A}^{\rm neq}\frac{d}{dt}\tilde{x}+\frac{d}{dt}\Sigma_{\rm A}$ $U_{\rm A}^{\rm neq},\tilde{x}$ $\Gamma_{\rm A}^{\rm neq}=\Lambda_{\rm A}^{\rm neq}+\tilde{x}{X}_{\rm A}^{\rm neq}$
II.2.2 Non-equilibrium Helmholtz–Gibbs potentials
Table 2: Time derivative forms of the non-equilibrium (neq) Helmholtz–Gibbs potentials as functions of intensive variables $T(t)$ and $x(t)$ and extensive variables $S_{\rm A}^{\rm neq}(t)$ and $X_{\rm A}^{\rm neq}(t)$, which are interrelated through the Legendre transformations shown in the final column. Because heat is always lost in non-equilibrium processes, the waste heat $Q_{\rm A}^{\rm wst}$ appears in the equations.
@@ II.2.2 Non-equilibrium Helmholtz–Gibbs potentials
Helmholtz Energy $\frac{d}{dt}F^{\rm neq}_{\rm A}=-S^{\rm neq}_{\rm A}\frac{d}{dt}T+x\frac{d}{dt}X_{\rm A}^{\rm neq}+\frac{d}{dt}Q_{\rm A}^{\rm wst}$ $T,x$
Gibbs Energy $\frac{d}{dt}G^{\rm neq}_{\rm A}=-S^{\rm neq}_{\rm A}\frac{d}{dt}T-X_{\rm A}^{\rm neq}\frac{d}{dt}x+\frac{d}{dt}Q_{\rm A}^{\rm wst}$ $T,X_{\rm A}^{\rm neq}$ $G^{\rm neq}_{\rm A}=F^{\rm neq}_{\rm A}-xX_{\rm A}^{\rm neq}$
Internal Energy $\frac{d}{dt}U^{\rm neq}_{\rm A}=T\frac{d}{dt}S^{\rm neq}_{\rm A}+x\frac{d}{dt}X_{\rm A}^{\rm neq}+\frac{d}{dt}Q_{\rm A}^{\rm wst}$ $S_{\rm A}^{\rm neq},x$ $U^{\rm neq}_{\rm A}=F^{\rm neq}_{\rm A}+TS^{\rm neq}_{\rm A}$
Enthalpy $\frac{d}{dt}H^{\rm neq}_{\rm A}=T\frac{d}{dt}S^{\rm neq}_{\rm A}-X_{\rm A}^{\rm neq}\frac{d}{dt}x+\frac{d}{dt}Q_{\rm A}^{\rm wst}$ $S_{\rm A}^{\rm neq},X_{\rm A}^{\rm neq}$ $H^{\rm neq}_{\rm A}=U^{\rm neq}_{\rm A}-xX_{\rm A}^{\rm neq}$
We introduce non-equilibrium Helmholtz and Gibbs energies defined as $F^{\rm neq}_{\rm A}(t)=-\Phi_{\rm A}^{\rm neq}(t)/\beta(t)$ and $G^{\rm neq}_{\rm A}(t)=-\Xi_{\rm A}^{\rm neq}(t)/\beta(t)$. Because $d\beta(t)/dt=-(1/k_{\rm B}T^{2}(t))dT(t)/dt$, we obtain the time derivative forms of these from Eqs. (14) and (15) as
@@
$\frac{dG_{\rm A}^{\rm neq}(t)}{dt}=-S_{\rm A}^{\rm neq}(t)\frac{dT(t)}{dt}-X_{\rm A}^{\rm neq}(t)\frac{dx(t)}{dt}+\frac{dQ_{\rm A}^{\rm wst}(t)}{dt},$ (20)
where $S_{\rm A}^{\rm neq}(t)=k_{\rm B}\Lambda_{\rm A}^{\rm neq}(t)$ is a non-equilibrium entropy, and we have introduced the waste heat current as
$\frac{dQ_{\rm A}^{\rm wst}(t)}{dt}=-\frac{1}{\beta(t)}\frac{d\Sigma_{\rm A}(t)}{dt}.$ (21)
From Eq. (16), we obtain the TDL transformation between the non-equilibrium Gibbs and Helmholtz potentials as
@@
Solving Eq. (17) for $U_{\rm A}^{\rm neq}(t)$, yields the following TDL transformations:
@@
From the above equations, we obtain the time derivative expressions for the enthalpy and internal energy. The results are summarized in Table 2. These results are reduced to those in Table 6 (Appendix A) for the quasi-static case. Thus, we can regard these potentials as extensions of thermodynamic potentials to a non-equilibrium regime.
@@
is the extensive heat current. From Eqs. (18) and (25), we can evaluate the entropy production rate as
$\frac{d\Sigma_{\rm A}(t)}{dt}=\frac{1}{k_{\rm B}}\left\{\frac{dS_{\rm A}^{\rm neq}(t)}{dt}-\frac{1}{T(t)}\frac{dQ_{\rm A}^{ext}(t)}{dt}\right\}.$ (26)
In the isothermal case, by integrating both sides of the two equilibrium states over time $t$, we obtainSakamoto and Tanimura (2020); Koyanagi and Tanimura (2022a)
$\Sigma_{\rm A}(t)=\frac{1}{k_{\rm B}}\left\{\Delta S_{\rm A}^{\rm qst}-\frac{Q_{\rm A}^{ext}}{T}\right\}.$ (27)
The non-equilibrium Gibbs energy satisfies
@@
$H_{\rm A}(t)=-T^{2}(t)\frac{\partial}{\partial T(t)}\left(\frac{G^{\rm neq}_{\rm A}(t)}{T(t)}\right)_{x(t),\tilde{Q}^{\rm wst}}.$ (30)
Equation (30) extends the Gibbs–Helmholtz relation to a non-equilibrium regime.